What is the perimeter of a rectangle with a length of 6 4/5 inches and a width of 2 2/3 inches?
Answer by sachi(548)

length of 6 4/5=34/5 inches and
a width of 2 2/3=8/3 inches
so perimeter=2(l+b)=2*(34/5+8/3)=2*(34*3+5*8)/5*3=284/15=18 14/15